OpenStudy (anonymous):

Medal and Fan!! The Georgia Peach Company ships peaches in wooden boxes. An empty wooden box has a mass of 2.5 kilograms. The peaches have an average mass of 0.15 kilogram each. Part a) Write an equation that expresses m, the total mass of a box of peaches, in terms of p, the number of peaches in the box. Part b) Solve the equation if there are 250 peaches. Show all steps.

OpenStudy (anonymous):

@stamp

OpenStudy (anonymous):

@bohotness

OpenStudy (anonymous):

@KayleeNeedsHelp

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (anonymous):

m=2.5+(.15p) m=2.5+.15*250 m=2.5+37.5 m=40kg

OpenStudy (bohotness):

Mhmm

OpenStudy (bohotness):

Right

OpenStudy (anonymous):

Thanks so much! That really helped!

OpenStudy (anonymous):

Glad to!

OpenStudy (spicytaco1):

k12?